the sides of the triangle are in the ratio 3:2:4 if the perimeter of the triangle is 27cms find the length of each side
Answer by edjones(8007)

3+2+4=9
3/9 * 27 = 9
2/9 * 27 = 6
4/9 * 27 = 12
.
Ed